### v1

@@ Abstract
Fibrinogen is an abundant protein found in human blood plasma that plays crucial roles in various physiological processes. When nanoparticles (NPs) are introduced into a biological solution, layers of biomolecules form on their surface, creating a corona. Understanding how the structure of the protein evolves into the corona is essential for evaluating the safety and toxicity of nanotechnology. However, the influence of NP properties on protein conformation is not well understood. In this study, we propose a new method that addresses this issue by analyzing multi-component spectral data using Machine Learning (ML). We apply the method to fibrinogen at blood concentrations while interacting with hydrophobic carbon or hydrophilic silicon dioxide NPs, revealing striking differences in the temperature dependence of the protein structure between the two cases. Our unsupervised ML method a) does not suffer from the challenges associated with the curse of dimensionality, and b) simultaneously handles spectral data from various sources. The method provides a deeper understanding of the correlation between protein structure and NP interactions, which could support the development of nanomedical tools to treat various conditions.
@@ 1 Introduction
However, the way the NP surface influences the protein structural changes within the corona remains debated [13]. For instance, theory predicts that proteins near a hydrophobic interface should unfold and aggregate less than those in bulk, and that increasing temperature does not facilitate unfolding and reduces aggregation [14]. Conversely, on a hydrophilic NP, the prediction is that proteins adsorb while remaining folded, and as temperature increases, they unfold but do not desorb [15].
Here, combining different experimental techniques with Machine Learning (ML) analysis, we characterise Fib structural changes over temperature when exposed to hydrophobic carbon NPs (CNP) and hydrophilic silicon dioxide (SiO2) NPs. We use several spectroscopic techniques: UltraViolet Resonance Raman (UVRR) spectroscopy [16], Circular Dichroism (CD) [17], and UV absorption [18] spectroscopy.
@@
We analyze our high-dimensional multi-component data from multiple spectroscopic techniques using a novel unsupervised ML workflow designed to quantify and visualize protein conformational changes as temperature varies. The novelty of our proposed ML model rests on two key algorithms: the Wasserstein distance [23, 24] and t-distributed stochastic neighbor embedding (t-SNE) [25, 26, 27].
@@ 3 ML model
The t-SNE is a nonlinear dimensionality reduction technique that converts high-dimensional data into a probability distribution reflecting the similarities among data points. It is specifically designed to address the crowding problem, which refers to the risk of losing local structure and compressing distances when high-dimensional data is mapped to a lower-dimensional space. The t-SNE mitigates the crowding problem because the heavy tails of the Student-t distribution in the low-dimensional space allow for better separation of points that are moderately distant. When applied to complex, high-dimensional data where local relationships are crucial, t-SNE outperforms the faster PCA, which cannot handle When applied to complex, high-dimensional data where local relationships are crucial, t-SNE outperforms the faster PCA, which cannot effectively handle data from different sources due to its reliance on variance maximization. This maximization can be strongly affected by outliers, leading to misleading results.
In our metric computations, we consider as reference spectra those taken at $T_{0}=22^{\circ}$, which is the lowest temperature of our measurements. At $T_{0}$, Fib is in its native state. As a result, we compute the similarity distance between the composite spectrum at $T_{0}$ and the other spectra at temperatures $T$, separated by $\Delta T=4^{\circ}$. Accordingly, the Wasserstein distance quantifies the increase of the protein’s structural disorder with temperature. To emphasize the change, we normalize the Wasserstein distance with the two values at the two extreme temperatures (Fig. 8).
We find that the disorder content of the protein increases monotonically with temperature for all systems. However, while the normalized disorder content for the free Fib case and the Fib with silica NPs depends on temperature in the same way, the case of Fib with carbon NPs stands out. The first two experiments clearly show that at around $(52\pm 2)^{\circ}$, the normalized Wasserstein distance experiences a significant change, while outside that range, the increase is approximately linear.
@@
In contrast, for the Fib with carbon NPs, the disorder increases linearly across the entire range of explored temperatures, without any characteristic $T$ (Fig. 8, central panel). Therefore, the impact of the NP surface is substantial, inducing only a gradual structural change. This analysis allows us to interpret the original spectra better, clarifying that the structural disordering induced by the increase of temperature in this case does not resemble the standard unfolding occurring for the free protein.
It is important to recall that we normalize the Wasserstein distance relative to the minimum and maximum disorder in each case. Therefore, this quantity does not inform us about the absolute value of disorder as a function of $T$ in each case. [^2] The absolute quantity is estimated in our next workflow step and varies in each instance. However, it is intriguing to observe that the $T$-dependence of the normalized quantity calculated for free Fib and Fib on silica NPs is remarkably similar regarding relative changes. In contrast, the differences with the CNP case suggest that hydrophilic silica and hydrophobic carbon affect Fib’s structure in fundamentally different ways.
@@
Figure 8: Normalized Wasserstein distance computed for composite spectra of free Fib and Fib with NPs. Upper panel: Fib (squares). Central panel: Fib with CNPs (circles). Lowe panel: Fib with SiO2NP (triangles). In each case, the quantity is normalized with respect to its values at $22^{\circ}$ and $86^{\circ}$. Therefore, it quantifies the relative distance of the spectra (structural disorder) from the reference spectrum of an ordered Fib protein (without or with NP) at $22^{\circ}$.
Next, to evaluate the absolute amount of disorder the Fib acquires with temperature increases in each case, we apply t-SNE to the data under scrutiny. This manifold learning algorithm models similarities between composite spectra in the original high-dimensional space using a Gaussian distribution and maps it into a Student-t distribution in the low-dimensional space [26]. It then minimizes the Kullback–Leibler (KL) divergence between the probability distribution of the original high-dimensional data and the low-dimensional space, which in our case is two, visualizing a corresponding two-dimensional embedding [26]. The KL divergence is a nonsymmetric measure of similarity between two distributions that, when minimized, preserves the local structure of the data. Hence, if two points are close in high-dimensional space, a minimum KL divergence keeps them close in the low-dimensional space as well.
To compute t-SNE, an input parameter called the perplexity $\mathcal{P}$ is required. The perplexity roughly corresponds to the effective number of neighbors in the high-dimensional space for each point considered when calculating similarities. In cases where several thousand data sets are available, the suggested default values for $\mathcal{P}$ are $30-50$ [27]. However, in our case, due to the small size of each dataset with only 17 temperatures, we propose to automatically find the optimal value of $\mathcal{P}$ through a grid search based on K-means clustering, a popular unsupervised machine learning algorithm used for clustering data, and the average silhouette score, which gives an overall measure of how appropriately the data has been clustered, as discussed in Section $4$ of ESI [55, 57]. According to our procedure, the perplexity is set to $7$, consistent with the small size of the datasets. Other parameters are set to their default values for t-SNE optimization [27].
@@
Figure 9: t-SNE analysis of composite spectra. a) Two-dimensional embedding from t-SNE with cosine similarity of composite spectra of Fib (squares), Fib with CNPs (circles), and Fib with SiO2NPs (triangles), each corresponding to a given temperature. Temperatures are indicated in the color bar, with $\Delta T=4^{\circ}$ starting from $T=22^{\circ}$. We used $\mathcal{P}=7$ and PCA initialization for t-SNE computations. b) Same as in panel a), with centroids (red cross marks) computed using the K-means clustering algorithm, assuming the optimal number of clusters $n_{\rm op}=5$ as input. In this case, the average silhouette score is $\approx 0.94$.
@@
To calculate the embedding, we compute t-SNE using cosine similarity (Eq. 1), as it is expected to make t-SNE less susceptible to the high dimensionality of the spectral data. We find five distinct clusters (Figure 9 a). All data corresponding to Fib with CNPs (circles) are grouped into a single cluster. The remaining data—Fib (squares) and Fib with SiO2NPs (triangles)—are divided into two clusters, each characterized by a distinct temperature range: $T\leq 50^{\circ}$ and $T\geq 54^{\circ}$.
@@
Therefore, the emergence of two clusters for Fib and Fib with SiO2NPs indicates a significant conformational change around $(52\pm 2)^{\circ}$, consistent with our normalized Wasserstein metrics. Furthermore, the calculation of centroids based on the K-means algorithm, with the number of clusters set to $5$, confirms that the data instances are well clustered. The clustering pattern has an average silhouette score of approximately $0.94$, indicating strong clustering [55] (Fig. 9 b).
Additionally, this analysis clearly demonstrates that the change for Fib with silica NPs is much smaller, approximately 20%, compared to free Fib. Therefore, although the Fib adsorbed onto hydrophilic silica NPs unfolds at the same temperature as the free protein, its structural change with increasing $T$ is about 20% smaller than in the free case.
Notably, the clustering into a single group for the Fib with CNPs composite spectra indicates that at all analyzed temperatures, no significant structural change occurs, again consistent with our normalized Wasserstein analysis. Therefore, the combination of t-SNE and normalized Wasserstein distance calculations emphasizes the gradual increase of a mildly structurally disordered state as a function of $T$ for Fib adsorbed on the hydrophobic carbon NPs.
@@
We emphasize that if PCA is used instead of t-SNE, the resulting two-dimensional projection would suggest the presence of five or more clusters. Furthermore, the calculation of the centroids would be less clear, making it more difficult to quantitatively evaluate the absolute amount of structural disorder in the three cases. This is because PCA displays some data points as outliers, as discussed in Section 3 of the ESI, due to PCA’s difficulty in appropriately handling data originating from very different sources.
@@
Finally, we note that a similar embedding is obtained when computing t-SNE with the Euclidean norm/distance and $\mathcal{P}=7$, as discussed in Section 4 of the ESI. In this case, the clustering quality decreases, with the average silhouette score dropping to approximately $0.91$. Five well-defined clusters still emerge, reinforcing the idea that t-SNE is not significantly affected by the data’s high dimensionality. However, the absolute disorder increase for the Fib with silica NPs would now be 43% of the one for free Fib, revealing that this estimate depends on the quality of the clustering.
@@ 4 Conlcusion
We developed a method that leverages ML to interpret spectral data from multiple experimental techniques measuring the structural properties of proteins. This approach optimizes the combined information and systematically evaluates the relative and absolute structural disorder of proteins as a function of varying external conditions.
@@
As a relevant case study, we consider the debated behavior of the effect of NP’s surface chemistry on the structural changes of proteins adsorbed within the corona at varying temperatures. To this goal, we consider fibrinogen at blood concentrations while interacting with hydrophobic carbon or hydrophilic silicon dioxide NPs. Our findings demonstrated that the NP chemistry has a significant impact on the stability of Fib as a function of $T$, with striking differences between the hydrophilic and the hydrophobic case.
@@
Specifically, our ML method demonstrates that the silica NPs reduce the Fib’s structural changes at increasing $T$ by 20% compared to the free Fib case. Furthermore, our approach identifies the adsorbed Fib unfolding transition temperature at $(52\pm 2)^{\circ}$, which coincides with the temperature at which free Fib unfolds. Above and below this characteristic temperature, the relative disorder of Fib adsorbed on the silica NPs follows the same linear trend as the free Fib. At high $T$, this trend is associated with an aggregation process. Therefore, the interaction with the hydrophilic surface moderates the protein unfolding but does not qualitatively change its aggregation compared to the free protein case. These experimental results confirm recent theoretical predictions based on simple modeling of hydrated proteins adsorbing onto a hydrophilic interface [15].
@@
The ML method shows that results are drastically different when Fib interacts with the hydrophobic carbon NPs. Our analysis reveals that the proteins interacting with CNPs do not undergo an unfolding transition at any of the explored temperatures between 22 and 86C, contrary to the bulk and silica NPs case. The Fib structure, instead, reaches a mildly disordered state that is stabilized against $T$ increases when adsorbed onto the hydrophobic CNP surface. This effect, in turn, reduces the aggregation phenomenon with temperature. Therefore, the ML method allows us to clearly confirm for the first time the stabilization effect of a hydrophobic surface against protein conformational changes and aggregation, as predicted by recent theoretical modeling [14].
@@
Overall, our ML workflow presents a unique perspective on the potential applications of protein corona by using NPs with specific properties. This is particularly significant given the increasing use of NPs in various medical contexts, such as drug delivery and imaging. These findings indicate that engineering NP surface properties can create new methods for modulating protein stability and may hold substantial promise in therapeutic and diagnostic applications. For example, adjusting the NPs’ surface to be more hydrophobic or hydrophilic could facilitate differential control over protein aggregation, thereby enhancing stability under physiological conditions.
@@
In conclusion, our novel use of ML algorithms provides a unique analytical perspective on the characterization of the phenomena. Future applications of these methods could explore broader environmental conditions such as pH, ionic strength, or the presence of other biomolecules in more complex biological milieus, further advancing the understanding of protein-NP interactions and their practical implications.

### v2

@@ Abstract
When nanoparticles (NPs) are introduced into a biological solution, layers of biomolecules form on their surface, creating a corona. Understanding how the protein’s structure evolves into the corona is essential for evaluating the safety and toxicity of nanotechnology. However, the influence of NP properties on protein conformation is not well understood. In this study, we propose a new method that addresses this issue by analyzing multi-component spectral data using machine learning (ML). We apply the method to fibrinogen, a crucial protein in human blood plasma, at physiological concentrations while interacting with hydrophobic carbon or hydrophilic silicon dioxide NPs, revealing striking differences in the temperature dependence of the protein structure between the two cases. Our unsupervised ML method a) does not suffer from the challenges associated with the curse of dimensionality, and b) simultaneously handles spectral data from various sources. The method offers a quantitative analysis of protein structural changes upon adsorption and enhances the understanding of the correlation between protein structure and NP interactions, which could support the development of nanomedical tools to treat various conditions.
@@ 1 Introduction
However, the way the NP surface influences the protein structural changes within the corona remains debated [13, 14, 15, 16]. For instance, theory predicts that proteins near a hydrophobic interface should unfold and aggregate less than those in bulk, and that increasing temperature does not facilitate unfolding and reduces aggregation [17]. Conversely, on a hydrophilic NP, the prediction is that proteins adsorb while remaining folded, and as temperature increases, they unfold but do not desorb [18].
Here, by combining different experimental techniques with Machine Learning (ML) analysis, we characterize the structural changes in Fib over temperature when exposed to hydrophobic carbon NPs (CNP) and hydrophilic silicon dioxide (SiO2) NPs. We employ several spectroscopic techniques: UltraViolet Resonance Raman (UVRR) spectroscopy [19], Circular Dichroism (CD) [20], and UV absorption [21] spectroscopy.
@@
We analyze our high-dimensional, multi-component data from various spectroscopic techniques using a novel unsupervised ML workflow designed to quantify and visualize how protein conformational changes occur as temperature varies. The novelty of our proposed ML model rests on two key algorithms: the Wasserstein distance [26, 27] and t-distributed stochastic neighbor embedding (t-SNE) [28, 29, 30].
@@ 3 ML model
The t-SNE is a nonlinear dimensionality reduction technique that converts high-dimensional data into a probability distribution reflecting the similarities among data points. It is specifically designed to address the crowding problem, which refers to the risk of losing local structure and compressing distances when high-dimensional data is mapped to a lower-dimensional space. The t-SNE mitigates the crowding problem because the heavy tails of the Student-t distribution in the low-dimensional space allow for better separation of points that are moderately distant. When applied to complex, high-dimensional data where local relationships are crucial, t-SNE outperforms the faster PCA, which cannot effectively handle data from different sources due to its reliance on variance maximization. This maximization can be significantly influenced by outliers, resulting in misleading results.
In our metric computations, we consider as reference spectra those taken at $T_{0}=22^{\circ}$, which is the lowest temperature of our measurements. At $T_{0}$, Fib is in its native state. As a result, we compute the similarity distance between the composite spectrum at $T_{0}$ and the other spectra at temperatures $T$, separated by $\Delta T=4^{\circ}$. Accordingly, the Wasserstein distance quantifies the increase of the protein’s structural disorder with temperature. To emphasize the change, we normalize the Wasserstein distance with the two values at the two extreme temperatures for the free Fib case (Fig. 8).
We find that the disorder content of the protein increases monotonically with temperature for all systems. In particular, the normalized disorder content for the free Fib and the Fib with silica NPs depends on temperature in a similar way. In both cases, around $(52\pm 2)^{\circ}$, the normalized Wasserstein distance experiences a significant change, while outside that range, the increase is approximately linear.
@@
However, while the free Fib attains the reference value of one at $T=86^{\circ}$, the protein with the silica NP only reaches a value of 0.9 at the same $T$. Hence, the silica surface partially limits the unfolding of Fib.
In contrast, the effect of the carbon interface is significantly different. In this case, the increase in the normalized Wasserstein distance is approximately linear across the entire range of explored temperatures, with no indication of a rapid structural change as in a folding-unfolding transition. Furthermore, the distance at $86^{\circ}$ only reaches a value of 0.5. Conventionally, a value of 0.5 for a folding-unfolding structural parameter is the limit between the two states. Furthermore, for the other two cases, the larger structural change occurs when the Wasserstein parameter is approximately 0.5.
Therefore, the impact of the NP surface is substantial, inducing only a gradual structural change. This analysis enables us to better interpret the original spectra, clarifying that the structural disorder induced by the increase in temperature in this case does not resemble the standard unfolding that occurs for the free protein.
These observations regarding the differences in the Wasserstein distances of Fib proteins interacting with hydrophilic silica NPs and hydrophobic carbon NPs indicate that the two interfaces impact Fib’s structure in fundamentally different ways. Furthermore, recalling that we normalize the Wasserstein distance relative to the minimum and maximum disorder in the free Fib case, Fig. 8 provides information about the absolute value of disorder as a function of $T$ in each case. In particular, Fib on silica NPs unfolds approximately 90% of the free case, while Fib on carbon NPs is about 50% folded at the same $T$.
@@
Figure 8: Normalized Wasserstein distance computed for composite spectra of free Fib and Fib with NPs. Upper panel: free Fib (squares). Central panel: Fib with CNPs (circles). Lowe panel: Fib with SiO2NP (triangles). In each case, the quantity is normalized with respect to the values at $22^{\circ}$ and $86^{\circ}$ for the free case, corresponding to Fib in its native (folded) or disordered (unfolded) state. Therefore, it quantifies the relative distance of the spectra (structural disorder) from the reference spectrum of a free Fib protein in its native state at $22^{\circ}$.
To better interpret the absolute amount of disorder the Fib acquires with temperature increase in each case, we apply t-SNE to the data under scrutiny. This manifold learning algorithm models similarities between composite spectra in the original high-dimensional space using a Gaussian distribution and maps it into a Student-t distribution in the low-dimensional space [29]. It then minimizes the Kullback–Leibler (KL) divergence between the probability distribution of the original high-dimensional data and the low-dimensional space, which in our case is two, visualizing a corresponding two-dimensional embedding [29]. The KL divergence is a non-symmetric measure of similarity between two distributions, which, when minimized, preserves the local structure of the data. Hence, if two points are close in high-dimensional space, a minimum KL divergence keeps them close in the low-dimensional space as well.
To compute t-SNE, an input parameter called the perplexity $\mathcal{P}$ is required. The perplexity roughly corresponds to the effective number of neighbors in the high-dimensional space for each point considered when calculating similarities. In cases where several thousand data sets are available, the suggested default values for $\mathcal{P}$ are between 30 and 50 [30]. However, in our case, due to the small size of each dataset with only 17 temperatures, we propose to automatically find the optimal value of $\mathcal{P}$ through a grid search based on K-means clustering, a popular unsupervised ML algorithm used for clustering data, and the average silhouette score, which gives an overall measure of how appropriately the data has been clustered, as discussed in Section 4 of ESI [58, 60]. According to our procedure, the perplexity is set to 7, consistent with the small size of the datasets. Other parameters are set to their default values for t-SNE optimization [30].
@@
Figure 9: t-SNE analysis of composite spectra. a) Two-dimensional embedding from t-SNE with cosine similarity of composite spectra of free Fib (squares), Fib with CNPs (circles), and Fib with SiO2NPs (triangles), each corresponding to a given temperature. Temperatures are indicated in the color bar, with $\Delta T=4^{\circ}$ starting from $T=22^{\circ}$. We used $\mathcal{P}=7$ and PCA initialization for t-SNE computations. b) Same as in panel a), with centroids (red cross marks) computed using the K-means clustering algorithm, assuming the optimal number of clusters $n_{\rm op}=5$ as input. In this case, the average silhouette score is $\approx 0.94$.
@@
To calculate the embedding, we compute t-SNE using cosine similarity (Eq. 1), as it is expected to make t-SNE less susceptible to the high dimensionality of the spectral data. We identify five distinct clusters (Figure 9a).
All data corresponding to free Fib (squares) and Fib with SiO2NPs (triangles) are divided into two clusters, each characterized by distinct temperature ranges: $T\leq 50^{\circ}$ and $T\geq 54^{\circ}$. The emergence of two clusters for free Fib and Fib with SiO2NPs indicates a significant conformational change around $(52\pm 2)^{\circ}$, consistent with our interpretation of the normalized Wasserstein metrics as a measure of protein disorder, which signifies the folding-unfolding transition of Fib as the temperature increases.
The remaining data for Fib with CNPs (circles) are grouped into a single cluster. This indicates that, at all analyzed temperatures, no significant structural change occurs, further supporting our rationale that a normalized Wasserstein distance with a value of approximately 0.5 corresponds to a protein that is around 50% folded.
@@
The calculation of centroids based on the K-means algorithm, with the number of clusters set to 5, confirms that the data instances are well clustered. The clustering pattern has an average silhouette score of approximately 0.94, indicating strong clustering [58] (Fig. 9 b).
This analysis demonstrates that, for normalized Wasserstein distance up to 0.5, the change in the Fib structure is limited, as with CNPs. Above the 0.5 value, the Wasserstein parameter marks an unfolded state, as with free Fib and Fib with silica NPs.
Although not quantitative, the t-SNE analysis qualitatively confirms that the structural change for Fib with SiO2NPs is less significant than in the free Fib case, even though the Fib unfolds at the same temperature in both scenarios. In contrast, it highlights the gradual increase of disorder without unfolding for Fib adsorbed on the hydrophobic carbon NPs up to $86^{\circ}$.
@@
We emphasize that if PCA is used instead of t-SNE, the resulting two-dimensional projection would suggest the presence of five or more clusters. Furthermore, the calculation of the centroids would be less clear, making it more difficult to quantitatively evaluate the absolute amount of structural disorder in the three cases. This is because PCA identifies some data points as outliers, as discussed in Section 3 of the ESI, due to its difficulty in handling data originating from very different sources.
@@
Finally, we note that a similar embedding is obtained when computing t-SNE with the Euclidean norm/distance and $\mathcal{P}=7$, as discussed in Section 4 of the ESI. In this case, the clustering quality decreases, with the average silhouette score dropping to approximately $0.91$. Nevertheless, five well-defined clusters still emerge, reinforcing the idea that t-SNE is not significantly affected by the data’s high dimensionality.
@@ 4 Conlcusion
We developed a method that leverages ML to interpret spectral data from multiple experimental techniques, measuring the structural properties of proteins. This approach optimizes the combined information and systematically evaluates the relative and absolute structural disorder of proteins as a function of varying external conditions.
@@
As a relevant case study, we consider the debated effect of NP surface chemistry on the structural changes of proteins adsorbed within the corona at varying temperatures. To achieve this goal, we consider fibrinogen at blood concentrations when interacting with hydrophobic carbon or hydrophilic silicon dioxide nanoparticles. Our findings demonstrate that the NP chemistry has a significant impact on the stability of Fib as a function of temperature, with striking differences between the hydrophilic and hydrophobic cases.
@@
Specifically, our ML method demonstrates that silica NPs reduce the Fib’s structural changes at increasing temperatures by 10% compared to the free Fib case. Furthermore, our approach identifies the adsorbed Fib unfolding transition temperature at $(52\pm 2)^{\circ}$, which coincides with the temperature at which free Fib unfolds. Above and below this characteristic temperature, the relative disorder of Fib adsorbed on the silica NPs follows the same linear trend as the free Fib. At high $T$, this trend is associated with an aggregation process. Therefore, the interaction with the hydrophilic surface moderates the protein unfolding but does not qualitatively change its aggregation compared to the free protein case. These experimental results confirm recent theoretical predictions based on simple modeling of hydrated proteins adsorbing onto a hydrophilic interface [18].
@@
The ML method reveals that the results are drastically different when Fib interacts with hydrophobic carbon NPs. Our analysis shows that the proteins interacting with CNPs do not undergo an unfolding transition at any of the explored temperatures between 22 and 86C, contrary to the bulk and silica NP cases. The Fib structure, on the other hand, reaches a mildly disordered state, without undergoing an unfolding transition, even at the highest $T$ explored when adsorbed onto the hydrophobic CNP surface. This effect, in turn, reduces the aggregation phenomenon with temperature. Therefore, the ML method enables us to clearly confirm, for the first time, the stabilization effect of a hydrophobic surface against protein conformational changes and aggregation, as predicted by recent theoretical modeling [17].
@@
Overall, our ML workflow presents a unique perspective on the potential applications of protein corona by using NPs with specific properties. This is particularly significant given the increasing use of NPs in various medical contexts, such as drug delivery and imaging. These findings suggest that engineering NP surface properties can lead to novel methods for modulating protein stability and hold substantial promise for therapeutic and diagnostic applications. For example, adjusting the NPs’ surface to be more hydrophobic or hydrophilic could facilitate differential control over protein aggregation, thereby enhancing stability under physiological conditions.
@@
In conclusion, our novel application of ML algorithms provides a unique analytical perspective on characterizing protein structural changes upon adsorption to nanomaterials. Future applications of these methods could explore broader environmental conditions, such as pH, ionic strength, or the presence of other biomolecules, in more complex biological milieus, thereby further advancing the understanding of protein-NP interactions and their practical implications.